# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ B.6 Hypothesis testing and confidence intervals
$\begin{gathered}\lim_{p\to\infty}\frac{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(P_{j}\leq\alpha,\beta_{j}^{\star}=0\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\beta_{j}^{\star}=0\right)}=\lim_{p\to\infty}\frac{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\left|\frac{\hat{r}_{*,j}-\beta_{j}^{\star}}{\sqrt{\hat{\tau}_{*}}}\right|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right),\absolutevalue{\beta_{j}^{\star}}\leq\frac{\mu_{0}}{2}\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\absolutevalue{\beta_{j}^{\star}}\leq\frac{\mu_{0}}{2}\right)}\\ =\frac{\mathbb{P}\left(|\mathsf{Z}|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right),\absolutevalue{\mathsf{B}^{\star}}\leq\frac{\mu_{0}}{2}\right)}{\mathbb{P}\left(\absolutevalue{\mathsf{B}^{\star}}\leq\frac{\mu_{0}}{2}\right)}=\mathbb{P}\left(|\mathsf{Z}|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right)\right)=\alpha\end{gathered}$
by Theorem 5.2 and Proposition A.5. Using exchangeability of columns of $\mathbf{X}$
@@
To see $(b)$, note that by Theorem 5.2 and Proposition A.5, almost surely
@@
Remark B.14 (Asymptotic limit of TPR).
@@
$\lim_{p\to\infty}\mathsf{TPR}(p)$ $=\lim_{p\to\infty}\frac{\sum_{j=1}^{p}\mathbb{I}\left(P_{j}\leq\alpha,\left|\beta_{j}^{\star}\right|\geq\mu_{0}\right)}{\sum_{j=1}^{p}\mathbb{I}\left(\absolutevalue{\beta_{j}^{\star}}\geq\mu_{0}\right)}$
$=\lim_{p\to\infty}\frac{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\left|\frac{r_{*,j}}{\sqrt{\hat{\tau}_{*}}}\right|\geq\Phi^{-1}\left(1-\frac{\alpha}{2}\right),\left|{\beta_{j}^{\star}}\right|\geq\mu_{0}\right)}{\frac{1}{p}\sum_{j=1}^{p}\mathbb{I}\left(\absolutevalue{\beta_{j}^{\star}}\geq\mu_{0}\right)}$
@@
where we used in the second line Theorem 5.2 and Proposition A.5.
Appendix C PCR-Spectrum-Aware Debiasing
C.1 Pseudo-code for PCR-Spectrum-Aware Debiasing
@@
$\hat{\bm{\beta}}_{\mathsf{co}}\leftarrow\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}_{\mathsf{new}}^{\top}(\mathbf{y}_{\mathsf{new}}-\mathbf{X}_{\mathsf{new}}\hat{\bm{\beta}})$ (101)
@@ C.2 Asymptotic normality
where we used that $J$ is finite not growing with $p$ and basic moment property of entries of $\mathbf{O}$ (see e.g. [65, Proposition 2.5]). It follows from a straightforward application of Markov inequality and Borel-Cantelli lemma that almost surely
@@
Using $J$ is finite, we obtain that almost surely,
@@
$\mathbf{D}_{\mathsf{new}}^{\top}\bm{1}_{(N-J)\times 1}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{D}.$
@@
C.3 Finite or single coordinate inference
Corollary C.1.

### v6

@@ C.8 Hypothesis testing and confidence intervals
[…]
by Theorem C.2 and Proposition A.5. Using exchangeability of columns of $\mathbf{X}$
@@
To see $(b)$, note that by Theorem C.2 and Proposition A.5, almost surely
@@
Remark C.17 (Asymptotic limit of TPR).
@@
[…]
@@
where we used in the second line Theorem C.2 and Proposition A.5.
Appendix D Proofs for PCR-Spectrum-Aware Debiasing
D.1 Pseudo-code for PCR-Spectrum-Aware Debiasing
@@
[…]
@@ D.2 Asymptotic normality
where we used that $J$ is finite not growing with $p$ and basic moment property of entries of $\mathbf{O}$ (see e.g. [79, Proposition 2.5]). It follows from a straightforward application of Markov inequality and Borel-Cantelli lemma that almost surely
@@
Using $J$ is finite and the assumption that $\limsup_{p\to\infty}\max_{i\in\mathcal{J}}d_{i}^{-2}/p\to 0$, we obtain that almost surely,
@@
$\mathbf{D}_{\mathsf{new}}^{\top}\bm{1}_{(N-J)\times 1}\equiv\mathbf{d}_{\mathcal{J}^{c}}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{D}.$
@@
D.3 Finite or single coordinate inference
Corollary D.1.